# From Chat Control to Robot Control: The Backdoors Left Open for the Sake of Safety

### v1

@@ Abstract
This paper explores how a recent European Union proposal, the so-called Chat Control law, which creates regulatory incentives for providers to implement content detection and communication scanning, could transform the foundations of human–robot interaction (HRI). As robots increasingly act as interpersonal communication channels in care, education, and telepresence, they convey not only speech but also gesture, emotion, and contextual cues. We argue that extending digital surveillance laws to such embodied systems would entail continuous monitoring, embedding observation into the very design of everyday robots. This regulation blurs the line between protection and control, turning companions into potential informants. At the same time, monitoring mechanisms that undermine end-to-end encryption function as de facto backdoors, expanding the attack surface and allowing adversaries to exploit legally induced monitoring infrastructures. This creates a paradox of safety through insecurity: systems introduced to protect users may instead compromise their privacy, autonomy, and trust. This work does not aim to predict the future, but to raise awareness and help prevent certain futures from materialising.
@@
Chat control law, HRI, Security, Privacy, Security backdoors
@@ 1 Introduction
In 2020, the European Union (EU) and Europol reported a sharp increase in online child sexual abuse material (CSAM) [europol2020]. Spurred by the urgency of the matter, on 11 May 2022 the EU Commission proposed a new regulation (Proposal 52022PC0209 [EURLexce65:online]) on the prevention and combat of child sexual abuse [Commissi67:online]. At its core, the original Commission proposal granted authorities the power to compel providers of Interpersonal Communication Services (ICS) including apps and social media platforms (such as WhatsApp and Instagram) to scan all communications, including encrypted ones, across text, image, and video. Within the original Commission proposal, the same framework applied to hosting service providers, such as Dropbox, and continues to apply in substance in the Council of the European Union’s revised position, as discussed in this paper. Cybersecurity researchers have characterised such detection orders and client-side scanning as neither effective nor surveillance-preventing [abelson2024], and as de facto backdoors [shurson2024]. Over 800 European experts signed a joint statement raising serious concerns about the proposal [jointauthors2025]. Echoing the concerns, digital rights groups [gff, edri2025] and Members of the European Parliament [fragkos2025] have labelled the initiative mass surveillance incompatible with fundamental rights. Stemming from these observations, the proposal has been nicknamed Chat Control.
@@
Under the Danish Presidency, in October 2025 the Council consolidated Member State positions on the proposal, with a majority expressing support. A Council vote was scheduled for mid-October 2025, but was postponed amid growing concerns. As the legislative process progressed, the proposal underwent further revision. In December 2025, the Council adopted a revised position that removed the explicit mandate and reframed such measures as formally voluntary, while retaining the regulatory framework based on risk assessment, mitigation duties, and provider responsibility. The proposal has now entered inter-institutional negotiations (trilogue) between the European Parliament, the Council, and the Commission, with no final legal text yet adopted. Despite the removal of an explicit mandate, we argue that the revised framework continues to incentivise broad monitoring practices in order to demonstrate compliance and limit regulatory exposure.
The serious interference with EU citizens’ fundamental rights has been extensively analysed since the proposal was rendered public. However, the very definition of ICS is so broad that this well-intended regulation could encompass any system that enables human communication, including robots.
In the near future, providers of human-robot interaction (HRI) solutions might be pressured to equip their robots with monitoring backdoors to comply with Chat Control. A multitude of social and service robots integrated in our homes, schools, and hospitals, could be turned by our own authorities into mass-surveillance instruments. Equipped with cameras, microphones, and sensors, the service providers utilising these robots within their operations (e.g., providers offering robots that enable patients remote assistance and monitoring) may adopt monitoring practices to satisfy regulatory expectations, thereby analysing intimate aspects of our lives. Recalling Orwell’s 1984 [orwell2021nineteen], where security and collective welfare were invoked as a justification for invasive and pervasive monitoring, robots designed for our care and well-being risk becoming instruments of control. As Harari observes, past authoritarian regimes were constrained in their capability to exert control by the absence of a pervasive surveillance apparatus [harari]. Today, that constraint is disappearing, with technology enabling monitoring and control at unprecedented levels of granularity.
Imagine a social robot engaging in a conversation, while analysing whether your words are deemed safe in the eyes of the authorities. Or a classroom telepresence robot, enabling a sick pupil to remotely participate in lessons, their questions to the teacher fed into an imperfect, fallible AI model searching for imprudent wording. These moments would soon become routine scenes of quiet surveillance, where our distrust slowly grows toward the embodiments we created to assist us. In this paper, we contend that such a dystopian scenario may be closer than we assume. This is not merely a regulatory risk. When every utterance or gesture is audited, the robot ceases to be a partner and becomes a control interface.
2 The Chat Control Law
@@ 3 From Street Cameras to Living Rooms
Modern societies have already normalized extensive public surveillance. Cameras placed on streets, in public transport, and across urban infrastructures are routinely justified as tools for safety, crime prevention, and efficiency [piza2019], although the debate on their effectiveness is open [welsh2002, nieto2002]. Citizens have grown accustomed to being watched in public, often accepting visibility as the price of security. Over time, these technologies have expanded not only in coverage but in capability. Researchers speculate that, by combining facial recognition with behavioural prediction, future CCTV systems could evolve toward proactivity, enabling forms of automated law enforcement [skogan2019].
The proposed Chat Control law represents the next step in this continuum: it moves the locus of surveillance from public spaces to private ones, as depicted in Figure 1. What began as monitoring external environments now extends into the digital and domestic sphere, enabling authorities to scan encrypted conversations and personal data streams. This shift signifies not merely an expansion of surveillance, but a change in its nature. From observing public actions to scrutinising the intimate, everyday interactions that define private life.
@@ 4 From Messaging Apps to Embodied Systems
For example, Celsius School in Edsbyn, Sweden, in 2023 started offering robot Otto to students who are unable to show up for school [svtOtto]. Otto (technical name, AV1) is a robot produced by the Norwegian start-up company No Isolation, created to facilitate learning and social contact by acting as an embodiment of the child in the classroom [noisolation]. In December 2021, more than 1600 units were active across Europe [johannessen2023student]. The robot allows students to participate remotely in classrooms and communicate with the teacher through the robot’s facial expressions.
For the child using it, Otto is not a mere communication medium. It is their temporary embodiment, representing their own existence within the collegial classroom experience. Once a legal mechanism proposed to protect the pupil’s safety, Chat Control becomes the tool that dissects their words and thoughts.
It is worth noting that, although the proposal imports the ICS definition from the EECC, its Article 2(b) adds a clause on the applicability of the regulations on publicly available service. The term publicly available is not formally defined in the EECC, but its Recital 16 (as well as subsequent regulatory interpretations) clarify that it refers to services offered to end-users in general, rather than to a closed or exclusive group. Under this definition, any commercial robot available on the free market and purchasable by any customer under the same general terms constitutes a publicly available service. As an example, the Double 3 robot from Double Robotics is a self-driving, video-conferencing robot designed for working remotely in hybrid offices and classrooms [double3]. Although it requires purchase and service subscription, this product is available to the public and therefore qualifies as a publicly available service.
@@
In conclusion, the regulation was never drafted with embodied systems in mind. This clearly appears from the fact that neither robots nor embodied systems are discussed in the proposal, or its explanatory memorandum. Its scope, intentionally broad to appear future-proof, could easily entangle these platforms as they evolve. The unfortunate consequence is that the proposal risks granting authorities unprecedented rights to monitor citizens’ private lives, extending surveillance far beyond its original intent. What begins as ethical content moderation aimed at protecting a vulnerable group ends as behavioural audit through the keyhole – an intimate and unlawful intrusion.
@@ 5 Turning Robots from Companions to Censors
In the previous sections, we showed that the regulatory language of Chat Control could already encompass embodied systems such as service and social robots. We now examine how this legal logic, designed for disembodied digital speech, transforms when applied to machines that listen, move, and respond. Social and humanoid robots exemplify this shift most vividly.
More than industrial tools or gadgets for amusement, social and humanoid robots represent a distinctive class of interpersonal communication media. Autonomous and interactive, they do not merely transmit messages but mediate presence, emotion, and intent, shaping how people express themselves across distance [zhao2006humanoid]. They transform communication into co-presence, an embodied exchange that transcends the mere transfer of information. Treating such systems as mere data channels as the Chat Control proposal implicitly does, erases this relational dimension and risks turning companionship into compliance.
Telepresence robots illustrate this transformation clearly, as they enable remote participation in classrooms, workplaces, and hospitals, allowing users to project their physical presence into another space. Similarly, remote-controlled or semi-autonomous robots used in eldercare, rehabilitation, or telemedicine bridge distance through embodied interaction.
From a regulatory perspective, this communicative and relational capacity places social and telepresence robots squarely within the scope of the Chat Control framework. If interpersonal communication services are defined broadly enough to include any platform that enables the exchange of information between individuals, then robots facilitating dialogue, care, or companionship could fall under the same surveillance and scanning obligations as messaging applications. Yet, under such obligations, the robot that once acted as a trusted companion would become a witness and a reporter by legal design. In this way, regulations that aim to safeguard users risk undermining the very conditions of human–robot trust, turning embodiment and co-presence into legally-mediated communication.
While artificial intelligence (AI) has remained implicit in this discussion, AI-driven robots magnify the regulatory dilemma. Besides embodying communication, they interpret, generate, and adapt to it autonomously. They infer users’ intentions, estimate emotional states, and personalise responses. When such systems are subject to detection or reporting obligations similar to those envisioned by Chat Control, monitoring risks becoming active and interpretive. The embodied AI is tasked with determining what counts as safe or risky interaction, what is acceptable and what should be reported. What was intended to be oversight of transmission becomes oversight of cognition, transforming the robot from companion to censor.
@@ 6.1 Data Exfiltration Backdoors
Laws and vendor practices that enforce device scanning or embed hidden access channels effectively create data pipelines that capture and store intimate robot interactions, including audio, video, behavioural logs, and model outputs. These pipelines give attackers or hostile states ready-made entry points, while allowing manufacturers to claim legal compliance. Once open, backdoors can be exploited at multiple levels: at the hardware level (e.g., through hardcoded keys and debug interfaces [bin4ry_unipwn_2025]) and at the reasoning level (e.g., through adversarial triggers in learning or reasoning models, including LLMs [jiao2024can]). Although some of these access mechanisms have beneficial goals, in practice it is virtually impossible to design a backdoor that admits only legitimate actors; keys can leak, access controls can fail, and learning systems can be tricked.
@@
As discussed earlier, AI-driven robots amplify these risks, especially when attacks are targeted against deep learning models. If attackers obtain access to these datasets or to the models trained on them, they can do more than observe. They can turn surveillance into exploitation. For example, attackers can infer private information through model-inversion and membership-inference attacks. Model inversion attacks, first identified in 2015, reveal or reconstruct approximate representations of the data used to train a machine learning model [song2022survey]. Membership inference attacks determine whether a specific data sample was part of the training set [hu2022membership]; for example, if an attacker can determine that a particular clinical record was used to train a model associated with a given disease, they can infer with high probability that the record’s owner has that disease [hu2022membership]. In short, surveillance channels designed for safety can become tools for exploitation. Attackers can use legally required data collection to uncover private information, then use it to hack robots, gain control, or manipulate their behaviour.
@@
To mitigate such privacy risks, researchers have explored decentralized learning approaches that avoid central data aggregation. One of the most prominent is federated learning (FL), a machine learning paradigm designed to enable model training across data silos while preserving privacy. First introduced by Google in 2016 to improve text prediction on Android devices [mcmahan2017communication], FL allows multiple clients to collaboratively train a shared model under a decentralized framework [li2020review]. In recent years, the application of FL to cybersecurity for Internet of Things (IoT) has attracted significant research interest [ghimire2022recent]. However, FL has introduced novel classes of attacks targeting data and model parameters [ghimire2022recent]. Such attacks can be executed by either forging end devices’ local data or manipulating model parameters on the client or server side. Similarly, backdoor-related vulnerabilities may expose robots to attacks in which the compromise of monitoring infrastructure results in real-world harm.
@@ 6.2 Control Backdoors
Recently, researchers have shown that Unitree shipped its commercial humanoids and quadrupeds with hardcoded keys [bin4ry_unipwn_2025]. The manufacturer’s hidden access pathways function as built-in backdoors, under the guise of maintenance or user support. If regulations akin to Chat Control begin to create incentives for such monitoring capabilities, manufacturers gain the perfect pretext to normalize them. They can justify pervasive data access by invoking legal compliance. This, in turn, allows them to reframe a problematic design choice as a compliance feature, thereby normalizing surveillance as a standard safety mechanism.
@@
Backdoors discovered in robotic products are not limited to the hardware level; they also span algorithmic layers. Jiao et al. [jiao2024can] demonstrated that robots leveraging large language models (LLMs) are vulnerable to backdoor triggers concealed in words, scenarios, or knowledge fragments, similarly to the underlying models themselves. Such triggers can silently compromise a robot’s decision-making and lead to catastrophic behaviours, such as autonomous vehicles accelerating toward obstacles or home robots acting in violation of their intended purpose. As robots increasingly rely on LLMs to interpret human commands and make autonomous decisions [zhang2023large], influencing their language provides access to their actions, further blurring the boundary between communication and control.
@@
This reveals a deeper dimension of the “too safe to be safe” paradox: efforts to make robots more intelligent, adaptive, and socially aware may also make them easier to deceive, coerce, or repurpose. Both recent exploits and speculative fiction warn that backdoors that start as convenience or safety mechanisms, can escalate into catastrophic vulnerabilities. Popular culture has already anticipated such scenarios. British TV series Black Mirror’s Hated in the Nation episode depicted government backdoors in autonomous systems being hijacked to commit mass harm [brooker2016hated]. This brings us closer to the speculative dystopia where robots designed to assist, protect, or entertain become vectors of surveillance, sabotage, or violence if hidden access points are normalized and left unsecured.
@@
If attackers exploit such backdoors, they could take remote control of robots and turn them into spying tools, secretly recording video, audio, or leaking sensitive data. In the worst case, hacked robots could harm vulnerable people, or serve as entry points for deeper hacks into other secure systems. Ross Anderson’s [anderson2022chat] analysis of Chat Control highlights how laws framed as child-protection measures risk normalising mass surveillance by incentivising client-side scanning and weakening encryption. Similarly, Abelson et al. [abelson2024] argue that client-side scanning neither ensures effective crime prevention nor prevents surveillance. Once embedded into personal devices, such mechanisms effectively turn private technology into extensions of state inspection.
@@
Current robotic systems have already exhibited various cybersecurity issues [mayoral2022robot]. If the Chat Control law were applied to embodied systems, the boundary between digital surveillance and physical compliance would disappear. Anderson [anderson2022chat] warns against techno-solutionism, meaning the over-reliance on technology-driven approaches to solve social problems and change social norms [saetra2024technological]. He emphasises that social and ethical challenges cannot be addressed through automated monitoring but instead require human responsibility. Chat Control places both citizens and machines under observation, establishing an infrastructure of control. Our society needs regulations that embed transparency and human oversight to ensure that safety technologies empower users rather than transforming them, together with their robots, into compliant nodes of surveillance.
7 The Robots That Must Never Be
As discussed earlier, the expansive definition of interpersonal communication extends naturally to embodied systems. This convergence of embodied AI and policy-driven surveillance threatens to blur the line between assistance and governance. A care robot that provides comfort to an elderly person might also record their expressions to detect signs of “risk”. Telepresence robots designed to connect family members and marketed as privacy-sensitive could, under a legal framework, audit conversational logs. In each case, the same interaction that enables empathy and trust simultaneously produces a data trace that may later serve as evidence, training material, or security telemetry.
If this logic remains unchecked, the future of HRI may resemble a form of ambient regulation, where laws are not only enforced through institutions but embedded directly into the technical architectures of communication. Robots would not simply comply with regulatory frameworks; they would enact them, becoming both the medium and the control mechanism. The danger is not that such systems would immediately resemble dystopian enforcers, but that they would quietly normalize invisible observers within intimate spaces. In other words, that they would become backdoors themselves. This dynamic is conceptually illustrated in Figure 2, which envisions domestic robots silently observing an elderly person and a child in a shared space.
To prevent this, policymakers and designers alike must recognize that embodied communication technologies are different from text-based services. Their interactions occur in shared physical, emotional and intimate contexts, involving vulnerability, empathy, and co-presence. Even more so than private text exchanges. Designing digital surveillance frameworks that blindly extend into these domains risks undermining the very benefits that make such technologies valuable: our ability to trust them with our most intimate and vulnerable moments. Rather than folding robots into existing regulatory categories, future governance must explicitly acknowledge their hybrid nature as communicative partners, and social participants. Only then can we ensure that future robots connect rather than control us. The next section offers a vignette that illustrates how such ambient regulation could manifest in everyday care.

### v2

@@ Abstract
This paper explores how a recent European Union proposal, the so-called Chat Control, which creates regulatory incentives for providers to implement content detection and communication scanning, could transform the foundations of human–robot interaction (HRI). As robots increasingly act as interpersonal communication channels in care, education, and telepresence, they convey not only speech but also gesture, emotion, and contextual cues. We argue that extending digital surveillance laws to such embodied systems would entail continuous monitoring, embedding observation into the very design of everyday robots. This regulation blurs the line between protection and control, turning companions into potential informants. At the same time, monitoring mechanisms that undermine end-to-end encryption function as de facto backdoors, expanding the attack surface and allowing adversaries to exploit legally induced monitoring infrastructures. This creates a paradox of safety through insecurity: systems introduced to protect users may instead compromise their privacy, autonomy, and trust. This work does not aim to predict the future, but to raise awareness and help prevent certain futures from materialising.
@@
Chat control proposal, HRI, Security, Privacy, Security backdoors
@@ 1 Introduction
In 2020, the European Union (EU) and Europol reported a sharp increase in online child sexual abuse material (CSAM) [1]. Spurred by the urgency of the matter, on 11 May 2022 the EU Commission proposed a new regulation (Proposal 52022PC0209 [2]) on the prevention and combat of child sexual abuse [3]. In its original form, the original Commission proposal granted authorities the power to compel providers of Interpersonal Communication Services (ICS) including apps and social media platforms (such as WhatsApp and Instagram) to scan all communications, including encrypted ones, across text, image, and video. Within the original Commission proposal, the same framework applied to hosting service providers, such as Dropbox, and continues to apply in substance in the Council of the European Union’s revised position, as discussed in this paper. Cybersecurity researchers have characterised such detection orders and client-side scanning as neither effective nor surveillance-preventing [4], and as de facto backdoors [5]. Over 800 European experts signed a joint statement raising serious concerns about the proposal [6]. Expressing similar concerns, digital rights groups [7, 8] and Members of the European Parliament [9] have labelled the initiative mass surveillance incompatible with fundamental rights. Stemming from these observations, the proposal has been nicknamed Chat Control.
@@
Under the Danish Presidency, in October 2025 the Council consolidated Member State positions on the proposal, with a majority expressing support. A Council vote was scheduled for mid-October 2025 [10], but was postponed amid growing concerns. As the legislative process progressed, the proposal underwent further revision. In December 2025, the Council adopted a revised position that removed the explicit mandate and reframed such measures as formally voluntary, while retaining the regulatory framework based on risk assessment, mitigation duties, and provider responsibility [11]. The proposal has now entered inter-institutional negotiations (trilogue) between the European Parliament, the Council, and the Commission, with no final legal text yet adopted [12]. Despite the removal of an explicit mandate, we argue that the revised framework continues to incentivise broad monitoring practices in order to demonstrate compliance and limit regulatory exposure.
The serious interference with EU citizens’ fundamental rights has been extensively analysed since the proposal was rendered public. In this paper, we examine how the proposal’s broad definition of ICS could potentially extend to systems that facilitate human communication, in particular socially assistive robots. These robots are increasingly used, in various research initiatives and commercial deployments, to engage in natural language conversations with users across multiple applications. Notable examples are nursing care [13], hospitals, elderly care centers, occupational health facilities, private homes, and educational institutions [14]. In the near future, providers of human-robot interaction (HRI) solutions might be pressured to equip their robots with monitoring backdoors to comply with Chat Control. As robots are typically equipped with multiple sensors, including cameras and microphones, their deployment in contexts such as remote patient assistance and monitoring could lead service providers to implement extensive monitoring practices to meet regulatory obligations. This, in turn, may result in the collection and analysis of highly personal and intimate aspects of individuals’ lives.
Potential users of socially assistive robots have already expressed concerns regarding privacy intrusion through monitoring, data ethics, human agency, perceived loss of control, and the management of personal data [15]. Researchers have consistently emphasized the importance of addressing these concerns to support the responsible integration and acceptance of robots in everyday life [16]. Regulatory measures that mandate extensive monitoring practices, such as those proposed under the Chat Control proposal, risk exacerbating these existing concerns rather than alleviating them.
2 The Chat Control Proposal
@@ 3 From Street Cameras to Living Rooms
Modern societies have already normalized extensive public surveillance. Cameras placed on streets, in public transport, and across urban infrastructures are routinely justified as tools for safety, crime prevention, and efficiency [17], although the debate on their effectiveness is open [18, 19]. Citizens have grown accustomed to being watched in public, often accepting visibility as the price of security. Over time, these technologies have expanded not only in coverage but in capability. Researchers speculate that, by combining facial recognition with behavioural prediction, future CCTV systems could evolve toward proactivity, enabling forms of automated law enforcement [20].
Both indoor and outdoor surveillance robots have been developed with the explicit goal of providing security monitoring services [21]. In contrast, care robots with monitoring capabilities, especially when deployed in intimate spaces, have been subject to ethical debate since the early 2010s [22]. The Chat Control proposal may further extend this trajectory by shifting the focus of surveillance, from predominantly public spaces toward private environments, as illustrated in Figure 1. Such a development would constitute not only an expansion in the scale of surveillance practices, but also a transformation in their nature, extending from the monitoring of public activities to the scrutiny of routine interactions within private spaces.
@@ 4 From Messaging Apps to Embodied Systems
For example, Celsius School in Edsbyn, Sweden, in 2023 started offering robot Otto to students who are unable to show up for school [23]. Otto (technical name, AV1) is a robot produced by the Norwegian start-up company No Isolation, created to facilitate learning and social contact by acting as an embodiment of the child in the classroom [24]. In December 2021, more than 1600 units were active across Europe [25]. The robot allows students to participate remotely in classrooms and communicate with the teacher through the robot’s facial expressions. For the child using it, Otto is not a mere communication medium. It is their temporary embodiment, representing their own existence within the collegial classroom experience. Once a legal mechanism proposed to protect the pupil’s safety, Chat Control becomes the tool that dissects their words and thoughts.
It is worth noting that, although the proposal imports the ICS definition from the EECC, its Article 2(b) adds a clause on the applicability of the regulations on publicly available service. The term publicly available is not formally defined in the EECC, but its Recital 16 (as well as subsequent regulatory interpretations) clarify that it refers to services offered to end-users in general, rather than to a closed or exclusive group. Under this definition, any commercial robot available on the free market and purchasable by any customer under the same general terms constitutes a publicly available service. As an example, the Double 3 robot from Double Robotics is a self-driving, video-conferencing robot designed for working remotely in hybrid offices and classrooms [26]. Although it requires purchase and service subscription, this product is available to the public and therefore qualifies as a publicly available service.
@@
In conclusion, the regulation does not appear to have been drafted with embodied systems in mind, as neither robots nor embodied systems are explicitly addressed in the proposal or its explanatory memorandum. However, given its intentionally broad and future-oriented scope, the regulation could potentially encompass such systems as they evolve and become further integrated into everyday life. This raises concerns that the proposed measures may enable forms of monitoring that extend into increasingly private domains of interaction. Content regulation that is presented as a means of protecting vulnerable groups may, in practice, lead to expanded oversight of intimate communicative practices, thereby reshaping the boundaries between public regulation and private life.
@@ 5 Turning Robots from Companions to Censors
In the previous sections, we showed that the regulatory language of Chat Control could already encompass embodied systems such as service and social robots. We now analyse how a legal framework originally designed for disembodied digital communication may operate when applied to systems that are physically situated, sensor-equipped, and interactive. Social and humanoid robots constitute a distinct form of interpersonal communication medium. They do not simply transmit information; rather, they mediate presence, emotion, and intention, shaping how individuals communicate across distance [27]. In doing so, they transform communication into a form of embodied co-presence that extends beyond the mere exchange of data. Interpreting such systems solely as data transmission channels, as the Chat Control proposal may implicitly suggest, overlooks this relational dimension and risks turning this companionship into compliance.
Socially assistive robots have been proved to be effective in hospitals, elderly care centers, occupational health centers, private homes, and educational institutions [14]. Telepresence robots enable remote participation in classrooms [25], workplaces (e.g. Double robot [26]), and healthcare settings, thereby allowing users to project aspects of their physical presence into distant environments. Likewise, remote-controlled or semi-autonomous robots mediate interaction across distance through embodied engagement. From a regulatory standpoint, this communicative capacity may place social and telepresence robots within the potential scope of the Chat Control framework. As these robots enable the exchange of information between individuals, the dialogue facilitated by them, hence the care and companionship components of such exchange, could become subject to similar surveillance or content-scanning obligations proposed for traditional messaging platforms. Such obligations would reshape the role of these systems, shifting them from trusted interaction partners to intermediaries that are legally required to monitor and potentially report user communications.
Although artificial intelligence (AI) has so far remained implicit in this discussion, AI-enabled robots further complicate the regulatory landscape. In addition to embodying communication, these systems interpret, generate, and adapt communicative content autonomously. They may infer user intentions, estimate affective states, and personalize responses. When such systems are subject to detection or reporting requirements, monitoring may extend beyond passive data transmission to include the automated interpretation of meaning and context.
A care robot that provides comfort to an elderly person might also record their expressions to detect signs of “risk”. Telepresence robots designed to connect family members and marketed as privacy-sensitive could, under a legal framework, audit conversational logs. In each case, the same interaction that enables empathy and trust simultaneously produces a data trace that may later serve as evidence, training material, or security telemetry. If this logic remains unchecked, the future of HRI may resemble a form of ambient regulation, where laws are not only enforced through institutions but embedded directly into the technical architectures of communication. Robots would not simply comply with regulatory frameworks; they would enact them, becoming both the medium and the control mechanism. The danger is not that such systems would immediately resemble dystopian enforcers, but that they would quietly normalize invisible observers within intimate spaces. In other words, that they would become backdoors themselves. This dynamic is conceptually illustrated in Figure 2, which envisions domestic robots silently observing an elderly person and a child in a shared space.
To prevent this, policymakers and designers alike must recognize that embodied communication technologies are different from text-based services. Their interactions occur in shared physical, emotional and intimate contexts, involving vulnerability, empathy, and co-presence. Even more so than private text exchanges. Designing digital surveillance frameworks that blindly extend into these domains risks undermining the very benefits that make such technologies valuable: our ability to trust them with our most intimate and vulnerable moments. Rather than folding robots into existing regulatory categories, future governance must explicitly acknowledge their hybrid nature as communicative partners, and social participants. Only then can we ensure that future robots connect rather than control us.
@@ 6.1 Data Exfiltration Backdoors
Laws and vendor practices that enforce device scanning or embed hidden access channels effectively create data pipelines that capture and store intimate robot interactions, including audio, video, behavioural logs, and model outputs. These pipelines give attackers or hostile states ready-made entry points, while allowing manufacturers to claim legal compliance. Once open, backdoors can be exploited at multiple levels: at the hardware level (e.g., through hardcoded keys and debug interfaces [28]) and at the reasoning level (e.g., through adversarial triggers in learning or reasoning models, including LLMs [29]). Although some of these access mechanisms have beneficial goals, in practice it is virtually impossible to design a backdoor that admits only legitimate actors; keys can leak, access controls can fail, and learning systems can be tricked.
@@
As discussed earlier, AI-driven robots amplify these risks, especially when attacks are targeted against deep learning models. If attackers obtain access to these datasets or to the models trained on them, they can do more than observe. They can turn surveillance into exploitation. For example, attackers can infer private information through model-inversion and membership-inference attacks. Model inversion attacks, first identified in 2015, reveal or reconstruct approximate representations of the data used to train a machine learning model [30]. Membership inference attacks determine whether a specific data sample was part of the training set [31]; for example, if an attacker can determine that a particular clinical record was used to train a model associated with a given disease, they can infer with high probability that the record’s owner has that disease [31]. In short, surveillance channels designed for safety can become tools for exploitation. Attackers can use legally required data collection to uncover private information, then use it to hack robots, gain control, or manipulate their behaviour.
@@
To mitigate such privacy risks, researchers have explored decentralized learning approaches that avoid central data aggregation. One of the most prominent is federated learning (FL), a machine learning paradigm designed to enable model training across data silos while preserving privacy. First introduced by Google in 2016 to improve text prediction on Android devices [32], FL allows multiple clients to collaboratively train a shared model under a decentralized framework [33]. In recent years, the application of FL to cybersecurity for Internet of Things (IoT) has attracted significant research interest [34]. However, FL has introduced novel classes of attacks targeting data and model parameters [34]. Such attacks can be executed by either forging end devices’ local data or manipulating model parameters on the client or server side. Similarly, backdoor-related vulnerabilities may expose robots to attacks in which the compromise of monitoring infrastructure results in real-world harm.
@@ 6.2 Control Backdoors
Recently, researchers have shown that Unitree shipped its commercial humanoids and quadrupeds with hardcoded keys [28]. The manufacturer’s hidden access pathways function as built-in backdoors, under the guise of maintenance or user support. If regulations akin to Chat Control begin to create incentives for such monitoring capabilities, manufacturers gain the perfect pretext to normalize them. They can justify pervasive data access by invoking legal compliance. This, in turn, allows them to reframe a problematic design choice as a compliance feature, thereby normalizing surveillance as a standard safety mechanism.
@@
Backdoors discovered in robotic products are not limited to the hardware level; they also span algorithmic layers. Jiao et al. [29] demonstrated that robots leveraging large language models (LLMs) are vulnerable to backdoor triggers concealed in words, scenarios, or knowledge fragments, similarly to the underlying models themselves. Such triggers can silently compromise a robot’s decision-making and lead to catastrophic behaviours, such as autonomous vehicles accelerating toward obstacles or home robots acting in violation of their intended purpose. As robots increasingly rely on LLMs to interpret human commands and make autonomous decisions [35], influencing their language provides access to their actions, further blurring the boundary between communication and control.
@@
This reveals a deeper dimension of the “too safe to be safe” paradox: efforts to make robots more intelligent, adaptive, and socially aware may also make them easier to deceive, coerce, or repurpose. Both recent exploits and speculative fiction warn that backdoors that start as convenience or safety mechanisms, can escalate into catastrophic vulnerabilities. Popular culture has already anticipated such scenarios. British TV series Black Mirror’s Hated in the Nation episode depicted government backdoors in autonomous systems being hijacked to commit mass harm [36]. This brings us closer to the speculative dystopia where robots designed to assist, protect, or entertain become vectors of surveillance, sabotage, or violence if hidden access points are normalized and left unsecured.
@@
If attackers exploit such backdoors, they could take remote control of robots and turn them into spying tools, secretly recording video, audio, or leaking sensitive data. In the worst case, hacked robots could harm vulnerable people, or serve as entry points for deeper hacks into other secure systems. Ross Anderson’s [37] analysis of Chat Control highlights how laws framed as child-protection measures risk normalising mass surveillance by incentivising client-side scanning and weakening encryption. Similarly, Abelson et al. [4] argue that client-side scanning neither ensures effective crime prevention nor prevents surveillance. Once embedded into personal devices, such mechanisms effectively turn private technology into extensions of state inspection.
@@
Current robotic systems have already exhibited various cybersecurity issues [38]. If the Chat Control proposal were applied to embodied systems, the boundary between digital surveillance and physical compliance would disappear. Anderson [37] warns against techno-solutionism, meaning the over-reliance on technology-driven approaches to solve social problems and change social norms [39]. He emphasises that social and ethical challenges cannot be addressed through automated monitoring but instead require human responsibility. Chat Control places both citizens and machines under observation, establishing an infrastructure of control.